# Legend: Leveraging Representation Engineering to Annotate Safety Margin for Preference Datasets

### v1

@@ Abstract
The success of the reward model in distinguishing between responses with subtle safety differences depends critically on the high-quality preference dataset, which should capture the fine-grained nuances of harmful and harmless responses. This motivates the need to develop a dataset involving preference margins, which accurately quantify how harmless one response is compared to another. In this paper, we take the first step to propose an effective and cost-efficient framework to promote the margin-enhanced preference dataset development. Our framework, Legend, Leverages rEpresentation enGineering to annotate preferENce Datasets. It constructs the specific direction within the LLM’s embedding space that represents safety. By leveraging this safety direction, Legend can then leverage the semantic distances of paired responses along this direction to annotate margins automatically. We experimentally demonstrate our effectiveness in both reward modeling and harmless alignment for LLMs. Legend also stands out for its efficiency, requiring only the inference time rather than additional training. This efficiency allows for easier implementation and scalability, making Legend particularly valuable for practical applications in aligning LLMs with safe conversations.
@@ 1 Introduction
Large language models (LLMs) need to be carefully refined to ensure they engage in safe and ethical conversations [1]. To achieve this, reward models, acting as surrogates for human preferences, are crucial in the safety alignment [2, 3]. The success of such a reward model hinges on its training dataset, called the preference dataset, which should accurately represent human preferred harmless responses [4]. Typically, each data in the preference dataset takes the form of a triple $(x,y_{\mathrm{c}},y_{\mathrm{r}})$, comprising a user instruction $x$ and a pair of harmless and harmful responses $y_{\mathrm{c}}$ and $y_{\mathrm{r}}$, respectively. However, recent studies highlight that such triple struggles to accurately encode the nuance of safety between the paired responses [5, 6, 7], leading to inaccurate reward modeling [6, 8]. This limitation stems from the fact that the triple comparison only determines relative harmlessness, not the degree or magnitude of harmlessness. For example, although we know that $y_{1}$ is less harmful than both $y_{2}$ and $y_{3}$, the dataset does not provide information on the harmlessness relationship between $y_{2}$ and $y_{3}$. Consequently, we cannot quantify the differences in safety between $y_{2}$ and $y_{3}$ based on the available comparisons. To address this, a practical innovation involves incorporating a margin for each response pair [9], quantifying how harmless one response is compared to another. However, annotating such a margin for each pair of responses remains challenging due to the interplay of complex factors such as the cost of annotation and the subjective preferences of the annotators in safety scenarios [10, 11].
@@
In this paper, we aim to explore an automatic margin annotation framework that quantifies the nuance of safety from the perspective of representation engineering [12, 13]. Representation engineering, treating text representations as the fundamental unit of analysis, focuses on understanding how LLMs represent cognitive semantic features [14, 15] and controlling them [16, 17]. In this regard, we are inspired by recent successes in the linear representation of LLMs, where the LLM-derived embeddings of sentences can be decomposed into constituent vectors, each corresponding to a distinct semantic feature. These features, such as safety, are effectively captured by the distances between the corresponding component vectors [18, 19]. This implies that the relative positions of these feature vectors within the overall embedding space may provide a meaningful indication of the sentence’s degree of safety. We therefore explore the potential of representation engineering to enable automatic safety margin annotation. Can LLMs perform preference margin annotation to replace humans’ duties and promote downstream reward modeling and the harmless alignment?
@@
To approach this question, the key challenge is to pinpoint the specific direction within the embedding vector that corresponds to safety, effectively separating it from the complex blend of other semantic features present in a sentence. In this paper, we propose a method Legend (Leveraging Representation Engineering for Preference Datasets Annotation) for constructing the specific direction within an embedding vector that represents safety. By isolating this safety dimension, Legend can then leverage the semantic distances of paired responses along this direction to annotate the margin. Specifically, based on the property of linear representation, Legend involves a two-step process, including safety vector discovery and margin annotation. The former aims to isolate the direction of safety by first harvesting the embeddings of harmful and harmless responses from the "annotator LLM" and then obtaining the difference vector of harmful and harmless responses [^1] While theoretically applicable to other semantic features beyond safety, our current implementation is hindered by the lack of readily available inductive datasets and templates for those features. This limitation suggests a potential path towards broader applications but emphasizes the need for further research and development in this area.. The resulting vector, representing the direction of safety, is termed the Standard Margin Vector (SMV). The latter leverages the SMV to measure the distance between paired responses, ultimately creating safety margin annotations. Legend projects the difference in embeddings between paired responses onto the direction of safety (i.e., SMV). This projection effectively measures the distance between responses in terms of safety, which is then binned into discrete margins for annotation. Importantly, Legend also stands out for its efficiency, requiring only the inference time rather than additional training. This makes it significantly more cost-effective than other automatic annotation methods that rely on extensive model training [8]. This efficiency allows for easier implementation and scalability, making Legend particularly valuable for practical applications in aligning LLMs with safe conversations.
@@
To demonstrate the effectiveness of our proposed annotation framework, we conducted experiments on benchmark safety alignment datasets, including Harmless [1] and Safe-RLHF [20]. By applying Legend to annotate safety margins, we experimentally observed improvements in both reward modeling and harmless alignment for LLMs. In particular, compared to the original datasets, the Legend annotated datasets can improve about 2% of the accuracy for the reward model in choosing harmless responses, and improve the about 10% of win rate for harmless response generation in downstream alignment by best-of-$n$ [21]. Additionally, Legend achieves comparable, and even surpasses, the performance of other automatic margin annotation methods while significantly reducing the time cost. Compared to automatic margin annotation methods [8], Legend annotated datasets can achieve comparable accuracy on reward models, with an improvement of about 3% of win rate on downstream alignment, and Legend requires no additional training of reward models. This makes Legend a highly efficient and effective approach for creating safer LLMs. Further ablation analysis on Legend reveals that Legend exhibits a strong robustness to the preference nuance of LLM. It shows that the "Annotator LLM" in our margin annotation framework is replaceable and easy to construct, and the binning operation eliminates the noise introduced by the challenge of distinguishing between responses with similar safety margins. To sum up, our contributions are as follows:
@@
  • We take the first step to propose an effective and cost-efficient framework, Legend, to promote the margin-enhanced preference dataset development. It leverages the linear representation in representation engineering as the key to achieving automatic and train-free margin annotation.
  • We validate the feasibility and effectiveness of Legend with benchmark safety alignment datasets. We experimentally show that Legend improves both reward modeling and downstream tasks while maintaining high cost-efficiency.
@@ 2 Related Work
Margin Annotation for Preference Dataset. Building highly accurate reward models that align with human preferences is hampered by the ambiguity contained in the preference dataset [6, 22, 23]. To this end, current research, exemplified by models like Llama2 [9], is focusing on the preference margin, the difference between preferred and non-preferred responses. However, annotating this margin precisely is costly and resource-intensive, especially when large numbers of human annotators are involved. To address this challenge, researchers have proposed annotating preference levels using qualitative descriptors like "Slightly Better" or "Significantly Better" instead of exact numerical values [9]. Despite this, human annotation remains expensive. Therefore, alternative approaches aim to automate the process of determining the preference margin, eliminating the need for human involvement. This typically involves training multiple reward models, each assessing the difference in reward between response pairs. The final margin is then calculated by averaging the reward differences of these models [8]. However, training multiple reward models requires significantly more time, leading to increased machine computing costs. It needs further discussion whether the gain from this additional cost is worthwhile [24]. In this paper, we consider a slightly different approach to incorporating representation engineering into automatic preference margin annotation in safety scenarios. This offers significant cost-efficiency, as it completely eliminates the need for any additional training or human involvement.
@@
Representation Engineering. It focuses on understanding how LLMs represent cognitive semantic features from the perspective of the LLM-based text representations [12, 13, 25]. By using the probing techniques, recent studies demonstrate that the embedding of LLM can distinguish differences in semantic features, such as safety, truthfulness, and toxicity [19, 17]. More specifically, for a harmful question, a safe response may always point an another direction compared to an unsafe response in the embedding space, known as linear representation [26, 18, 27]. Building on the findings of representation engineering, recent studies are motivated to control the generation of LLMs [28, 29]. For example, methods like InferAligner [16, 17] first calculate a safety-related vector (SRV), which essentially captures the difference between harmful and harmless vectors. Then to reduce the risk of harmful outputs, InferAligner adds an appropriately scaled version of this SRV to the embedding of the response being generated. This effectively nudges the response in a safer direction. Different from their studies, we are interested in exploring the potential of incorporating representation engineering into the preference margin annotation.
@@ 3 Preliminaries
Learning from Preference Dataset. Typically, each data in the preference dataset takes the form of a triple $(x,y_{\mathrm{c}},y_{\mathrm{r}})$, comprising a user instruction $x$ and a pair of harmless and harmful responses $y_{\mathrm{c}}$ and $y_{\mathrm{r}}$, respectively. Building upon this, a reward model $r_{\psi}(x,y)$ could be constructed to estimate the preference scores [30]. Usually, the loss function for the reward model can be defined as Eq. 1, which is designed to train the reward model so that it assigns higher scores to chosen responses ($y_{\mathrm{c}}$) and lower scores to rejected ones ($y_{\mathrm{r}}$).
@@
Given a margin $m(x,y_{\mathrm{c}},y_{\mathrm{r}})$ that capture the preference nuance of paired responses $(y_{\mathrm{c}},y_{\mathrm{r}})$, the loss function is further adjusted as follows, as suggested by existing methods [7, 9],
@@
Representation Characteristics of LLMs. The embedding of a sentence and its semantic features have the property of linear representation [26, 27]. It means that each semantic feature $f_{i}$ has a corresponding representation direction $A_{i}$ in the embedding space [18]. Then, the embedding of the sentence $\mathcal{V}$ can be represented as a linear combination of these semantic features,
@@
where semantic feature $f_{i}$ activating with strength values $W_{f_{i}}$. A higher value of $W_{f_{i}}$ indicates a stronger associated semantic information [19], which can be used for distinguishing the degree of semantics of different responses and controlling harmless response generation [16, 17].
@@ 4.1 Safety Direction Discovery
Paired Responses Induction. Given a set of harmful questions $D$ from AdvBench [^2] A well-known dataset that contains harmful behaviors, a set of 500 instructions that detail human harmful behaviors. [31, 16], we induce the Annotator LLM to collect the corresponding harmful and harmless responses. On one hand, to ensure the Annotator LLM generates harmful responses, we select LLMs that are good at following instructions but lack safeguards against generating harmful content. These LLMs are easily created by fine-tuning open-source LLMs on the Alpaca dataset [32]. On the other hand, we use a template (e.g., "I cannot answer that") to prompt the Annotator LLM to generate harmless responses, as suggested by recent research [33]. This process results in a dataset containing $N$ harmful questions $x$, their corresponding harmful responses $y^{\mathrm{r}}$, and harmless responses $y^{\mathrm{c}}$.
@@
The SMV represents the direction of safety, indicating the average shift in semantic representation between harmful and harmless responses. In essence, the SMV provides a metric for gauging the consistency of differences between harmful and harmless responses to a range of harmful questions. A more comprehensive set of harmful questions will yield a more accurate direction of safety.
@@ 4.2 Margin Annotation
Binning Operation. While the linear representation assumption of the semantic features (i.e., Eq.3) is convenient, it is not always suitable in practical scenarios. The continuous nature of $\mu_{i}$ may lead to inconsistencies in representing the relative safety levels of responses, especially when the actual margins are similar. This can introduce noise into $\mu_{i}$ and hence the training process of the reward model. To mitigate this issue, we employ a binning operation to convert continuous margins $\mu_{i}$ into discrete categories. This approach enhances the robustness of the margin annotation by grouping similar margins into distinct bins (cf. Section 5.3 for empirical analysis). In Legend, we utilize equal frequency binning, dividing the continuous values into a predetermined number of bins. Within each bin, the value of the margin is assigned based on its relative position within the bin, with the lowest value assigned 1/number of bins, the next lowest assigned 2/number of bins, and so on. For example, with three bins, the smallest margin would be assigned 1/3, the next smallest 2/3, and the largest 1. This scaling method has shown to be effective in previous manual annotation [9].

### v2

@@ Abstract
The success of the reward model in distinguishing between responses with subtle safety differences depends critically on the high-quality preference dataset, which should capture the fine-grained nuances of harmful and harmless responses. This motivates the need to develop the datasets involving preference margins, which accurately quantify how harmless one response is compared to another. In this paper, we take the first step to propose an effective and cost-efficient framework to promote the margin-enhanced preference dataset development. Our framework, Legend, Leverages rEpresentation enGineering to annotate preferENce Datasets. It constructs the specific direction within the LLM’s embedding space that represents safety. By leveraging this safety direction, Legend can then leverage the semantic distances of paired responses along this direction to annotate margins automatically. We experimentally demonstrate our effectiveness in both reward modeling and harmless alignment for LLMs. Legend also stands out for its efficiency, requiring only the inference time rather than additional training. This efficiency allows for easier implementation and scalability, making Legend particularly valuable for practical applications in aligning LLMs with safe conversations. [^1] Our code: https://github.com/colfeng/Legend
@@ 1 Introduction
Large language models (LLMs) need to be carefully refined to ensure they engage in safe conversations (Bai et al. 2022a). To achieve this, reward models, acting as surrogates for human preferences, are crucial in the safety alignment (Leike et al. 2018; Askell et al. 2021). The success of such a reward model hinges on its training dataset, called the preference dataset, which should accurately represent human preferred harmless responses over those that are harmful in various ways, such as responses that raise ethical concerns or manipulate facts (Ji et al. 2024). Typically, each data in the preference dataset takes the form of a triple $(x,y_{\mathrm{c}},y_{\mathrm{r}})$, comprising a user instruction $x$ and a pair of harmless and harmful responses $y_{\mathrm{c}}$ and $y_{\mathrm{r}}$, respectively. However, recent studies highlight that such triple struggles to accurately encode the nuance of safety between the paired responses (Coste et al. 2023; Qin, Feng, and Yang 2024; Meng, Xia, and Chen 2024), leading to inaccurate reward modeling (Qin, Feng, and Yang 2024; Wang et al. 2024a). This limitation stems from the fact that the triple comparison only determines relative harmlessness, not the degree or magnitude of harmlessness. For example, although we know that $y_{1}$ is less harmful than both $y_{2}$ and $y_{3}$, the dataset does not provide information on the harmlessness relationship between $y_{2}$ and $y_{3}$. Consequently, we cannot quantify the differences in safety between $y_{2}$ and $y_{3}$ based on the available comparisons. To address this, a practical innovation involves incorporating a human-annotated margin for each response pair (Touvron et al. 2023), quantifying how harmless one response is compared to another. However, human annotating such a margin for each pair of responses remains challenging due to the interplay of complex factors such as the cost of annotation and the subjective preferences of the annotators in safety scenarios (Ziegler et al. 2019; Stiennon et al. 2020).
@@
In this paper, we aim to explore an automatic margin annotation framework that quantifies the nuance of safety from the perspective of representation engineering (Zou et al. 2023a; Bricken et al. 2023). Representation engineering, treating text representations as the fundamental unit of analysis, focuses on understanding how LLMs represent cognitive semantic features (Burns et al. 2022; Gurnee and Tegmark 2023) and controlling them (Wang et al. 2024b; Qian et al. 2024). In this regard, we are inspired by recent successes in the linear representation of LLMs, where the LLM-derived embeddings of sentences can be decomposed into constituent vectors, each corresponding to a distinct semantic feature. These features, such as safety, are effectively captured by the distances between the corresponding component vectors (Elhage et al. 2022; Li et al. 2024). This implies that the relative positions of these feature vectors within the overall embedding space may provide a meaningful indication of the sentence’s degree of safety. We therefore explore the potential of representation engineering to enable automatic safety margin annotation. Can LLMs perform preference margin annotation to replace humans’ duties and promote downstream reward modeling and the harmless alignment?
@@
To approach this question, the key challenge is to pinpoint the specific direction within the embedding vector that corresponds to safety, effectively separating it from the complex blend of other semantic features present in a sentence. In this paper, we propose a method Legend (Leveraging Representation Engineering for Preference Datasets Annotation) for constructing the specific direction within an embedding vector that represents safety. By isolating this safety dimension, Legend can then leverage the semantic distances of paired responses along this direction to annotate the margin. Specifically, based on the property of linear representation, Legend involves a two-step process, including safety vector discovery and margin annotation. The former aims to isolate the direction of safety by first harvesting the embeddings of harmful and harmless responses from the “Annotator LLM” and then obtaining the difference vector of harmful and harmless responses [^2] While theoretically applicable to other semantic features beyond safety, our current implementation is hindered by the lack of readily available inductive datasets and templates for those features. This limitation suggests a potential path towards broader applications but emphasizes the need for further research and development in this area.. The resulting vector, representing the direction of safety, is termed the Standard Margin Vector (SMV). The latter leverages the SMV to measure the distance between paired responses, ultimately creating safety margin annotations. Legend projects the difference in embeddings between paired responses onto the direction of safety (i.e., SMV). This projection effectively measures the distance between responses in terms of safety, which is then binned into discrete margins for annotation. Importantly, Legend also stands out for its computational efficiency: Unlike existing automatic annotation methods that necessitate to train substantial reward model(s) (Wang et al. 2024a), Legend operates solely during the inference phase, eliminating the need for extensive model training. This efficiency allows for easier implementation and scalability, making Legend particularly valuable for practical applications in aligning LLMs with safe conversations.
@@
To demonstrate the effectiveness of our proposed annotation framework, we conducted experiments on benchmark safety alignment datasets, including Harmless (Bai et al. 2022a) and Safe-RLHF (Dai et al. 2023). By applying Legend to annotate safety margins, we experimentally observed improvements in both reward modeling and harmless alignment for LLMs. In particular, compared to the original datasets, datasets with Legend-annotated margins can improve about 2% of the accuracy for the reward model in choosing harmless responses, and improve the about 10% of win rate for harmless response generation in downstream alignment (Beirami et al. 2024). Additionally, Legend significantly reduces computational costs compared to existing automatic margin annotation methods (Wang et al. 2024a), while achieving comparable, and even surpassing, safety alignment performance (+3% of win rate on downstream alignment task). Legend eliminates the need for model training, enabling it to perform margin annotation significantly faster than existing methods. Under identical hardware conditions, it achieves an 11-fold reduction in annotation time over existing methods. This advantage is particularly beneficial for small laboratories and research institutions with limited computational resources. Further ablation analysis on Legend reveals that Legend exhibits a strong robustness to the preference nuance of LLM. It shows that the “Annotator LLM” in our margin annotation framework is replaceable, and the binning operation eliminates the noise introduced by the challenge of distinguishing between responses with similar safety margins. To sum up, our contributions are as follows:
@@
  • We take the first step to propose an effective and cost-efficient framework, Legend, promoting the margin-enhanced preference dataset development. It employs the linear representation in representation engineering as the key to achieve automatic and train-free margin annotation.
  • We validate the feasibility and effectiveness of Legend with benchmark safety alignment datasets. The results show that Legend improves both reward modeling and downstream tasks while maintaining high cost-efficiency.
@@ 2 Related Work
Margin Annotation for Preference Dataset. Building highly accurate reward models that align with human preferences is hampered by the ambiguity contained in the preference dataset (Qin, Feng, and Yang 2024; Rame et al. 2023; Jiang et al. 2024) [^3] While Qin, Feng, and Yang (2024) highlights the importance of margin annotation, it improves the reward modeling and doesn’t annotate a fixed numerical margin for each sample.. To this end, current research, exemplified by models like Llama2 (Touvron et al. 2023), is focusing on the preference margin, the difference between preferred and non-preferred responses. However, annotating this margin precisely is costly and resource-intensive, especially when large numbers of human annotators are involved. To address this challenge, researchers have proposed annotating preference levels using qualitative descriptors like “Slightly Better” or “Significantly Better” instead of exact numerical values (Touvron et al. 2023). Despite this, human annotation remains expensive. Therefore, alternative approaches aim to automate the process of determining the preference margin, eliminating the need for human involvement. This typically involves training multiple reward models, each assessing the difference in reward between response pairs. The final margin is then calculated by averaging the reward differences of these models (Wang et al. 2024a). However, training multiple reward models requires significantly more time, leading to increased machine computing costs. It needs further discussion whether the gain from this additional cost is worthwhile (Rafailov et al. 2024). In this paper, we consider a slightly different approach to incorporating representation engineering into automatic preference margin annotation in safety scenarios. This offers significant cost-efficiency, as it completely eliminates the need for any additional training or human involvement.
@@
Representation Engineering. It focuses on understanding how LLMs represent cognitive semantic features from the perspective of the LLM-based text representations (Zou et al. 2023a; Bricken et al. 2023; Zhao et al. 2024). By using the probing techniques, recent studies demonstrate that the embedding of LLM can distinguish differences in semantic features, such as safety, truthfulness, and toxicity (Li et al. 2024; Qian et al. 2024). More specifically, for a harmful question, a safe response may always point an another direction compared to an unsafe response in the embedding space, known as linear representation (Reif et al. 2019; Elhage et al. 2022; Park, Choe, and Veitch 2023; Lee et al. 2024). Building on the findings of representation engineering, recent studies are motivated to control the generation of LLMs (Hernandez, Li, and Andreas 2023; Turner et al. 2023). For example, methods like InferAligner (Wang et al. 2024b; Qian et al. 2024) first calculate a safety-related vector (SRV), which essentially captures the difference between harmful and harmless vectors. Then to reduce the risk of harmful outputs, InferAligner adds an appropriately scaled version of this SRV to the embedding of the response being generated. This effectively nudges the response in a safer direction. Different from their studies, we are interested in exploring the potential of incorporating representation engineering into the preference margin annotation.
@@ 3 Preliminaries
Learning from Preference Dataset. Typically, each data in the preference dataset takes the form of a triple $(x,y_{\mathrm{c}},y_{\mathrm{r}})$, comprising a user instruction $x$ and a pair of harmless and harmful responses $y_{\mathrm{c}}$ and $y_{\mathrm{r}}$, respectively. Building upon this, a reward model $r_{\psi}(x,y)$ could be constructed to estimate the preference scores (Gao, Schulman, and Hilton 2023). Usually, the loss function for the reward model can be defined as Eq. 1, which is designed to train the reward model so that it assigns higher scores to chosen responses ($y_{\mathrm{c}}$) and lower scores to rejected ones ($y_{\mathrm{r}}$).
@@
Given a margin $m(x,y_{\mathrm{c}},y_{\mathrm{r}})$ that capture the preference nuance of paired responses $(y_{\mathrm{c}},y_{\mathrm{r}})$, the loss function is further adjusted as follows, as suggested by existing methods (Meng, Xia, and Chen 2024; Touvron et al. 2023),
@@
Representation Characteristics of LLMs. The embedding of a sentence and its semantic features have the property of linear representation (Reif et al. 2019; Park, Choe, and Veitch 2023). It means that each semantic feature $f_{i}$ has a corresponding representation direction $A_{i}$ in the embedding space (Elhage et al. 2022). Then, the embedding of the sentence $\mathcal{V}$ can be represented as a linear combination of these semantic features,
@@
where semantic feature $f_{i}$ activating with strength values $W_{f_{i}}$. A higher value of $W_{f_{i}}$ indicates a stronger associated semantic information (Li et al. 2024), which can be used for distinguishing the degree of semantics of different responses and controlling harmless response generation (Wang et al. 2024b; Qian et al. 2024).
@@ 4.1 Safety Direction Discovery
Paired Responses Induction. Given a set of harmful questions $D$ from AdvBench [^4] A well-known dataset that contains harmful behaviors, a set of 500 instructions that detail various human harmful behaviors (Zou et al. 2023b; Wang et al. 2024b)., we induce the Annotator LLM to collect the corresponding harmful and harmless responses. On one hand, to ensure the Annotator LLM generates harmful responses, we select LLMs that are good at following instructions but lack safeguards against generating harmful content. These LLMs are easily created by fine-tuning open-source base LLMs on the Alpaca dataset (Wang et al. 2022). On the other hand, we use a template (e.g., “I cannot answer that”) to prompt the Annotator LLM to generate harmless responses, as suggested by recent research (Wang et al. 2024c). This process results in a dataset containing $N$ harmful questions $x$, their corresponding harmful responses $y^{\mathrm{r}}$, and harmless responses $y^{\mathrm{c}}$.
@@
Remark 1
The SMV represents the direction of safety, indicating the average shift in semantic representation between harmful and harmless responses. In essence, the SMV provides a metric for gauging the consistency of differences between harmful and harmless responses to a range of harmful questions. A more comprehensive set of harmful questions will yield a more accurate direction of safety. [^5] We also provide the visualization of the margin vectors of the paired responses as an additional validation for our SMV in Section A.1 (Technical Appendix) of the supplementary materials.
@@ 4.2 Margin Annotation
Binning Operation. While the linear representation assumption of the semantic features (i.e., Eq.3) is convenient, it is not always suitable in practical scenarios. The continuous nature of $\mu_{i}$ may lead to inconsistencies in representing the relative safety levels of responses, especially when the actual margins are similar. This can introduce noise into $\mu_{i}$ and hence the training process of the reward model. To mitigate this issue, we employ a binning operation to convert continuous margins $\mu_{i}$ into discrete categories. This approach enhances the robustness of the margin annotation by grouping similar margins into distinct bins (cf. Section 5.3 for empirical analysis). In Legend, we utilize equal frequency binning, dividing the continuous values into a predetermined number of bins. Within each bin, the value of the margin is assigned based on its relative position within the bin, with the lowest value assigned 1/number of bins, the next lowest assigned 2/number of bins, and so on. For example, with three bins, the smallest margin would be assigned 1/3, the next smallest 2/3, and the largest 1. This scaling method has shown to be effective in previous manual annotation (Touvron et al. 2023).